OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (anonymous):

a(b-c)=ab-ac

OpenStudy (anonymous):

Distribute: 10x-35=33-18x-12 Combine: 10x-35=21-18x 28x-35=21 28x=56 x=2

OpenStudy (anonymous):

10x-35=33-18x-12 10x-35=21-18x 10x+18x=21+35 28x=56 x=2